ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 40, 60 Butković, Davor, University of Zagreb, Croatia Transformations of regular almost convergent sequences The author studies the extension of the notion of regular convergence of multiple sequences in the sense of G. H. Hardy and F. Móricz to Cesàro convergent and almost convergent sequences. Various classes of almost convergent sequences are defined and for some of them necessary and sufficient conditions on matrices which transform one class to another are found. Also, rearrangements, dilutions, block decompositions and superpositions of double sequences are considered. There are applications to random chaines and random fields. Reference: D. Butković, On regular almost convergence, Mathematical Communications 1(1996), 127-137.